A piece of silver of mass 362 g has a heat capacity of 85.7 J/°C. What is the specific heat of silver?


1
Expert's answer
2021-09-22T00:26:42-0400

We calculate the specific heat capacity of silver by dividing the heat capacity of silver(Ag) by its mass;


"S_{Ag}=\frac{ C_{Ag}}{m_{Ag}}"


"=\frac{85.7J/^0C}{362g}"


"=0.237\frac{J}{g.^0C}"